# MIBench: Evaluating Multimodal Large Language Models over Multiple Images

### v1

@@ 4.2 Evaluation Results
Figure 3: A qualitative case of Subtle Difference task.
@@ 4.2.1 Multi-Image Instruction & Multimodal Knowledge-Seeking
Close-source MLLMs exhibit superior performance than close-source models. In most MII and MKS tasks, the performance of open-source models lags significantly behind that of proprietary models. For instance, on the Subtle Difference (SD) task, the best-performing open-source model Mantis achieves an accuracy of 54.1%, falling behind GPT-4o by 36.4%.
Open-source MLLMs are inadequate in fine-grained perception tasks. Although many open-source MLLMs have decent performance on the General Comparison (GC) task, their performance is significantly worse on the Subtle Difference (SD) task. For instance, IDEFICS2-I achieves 83.1% and 49.7% on the GC and SD tasks respectively. In contrast, GPT-4V and GPT-4o largely outperform open-source models, due to their high-resolution input strategy. Furthermore, Figure 3 provides a qualitative case supporting the above point.
Multi-image MLLMs perform better than single-image models on most tasks. This verifies that pre-training on interleaved image-text data (e.g. IDEFICS2) and instruction tuning on multi-image data (e.g. Mantis) are beneficial for improving multi-image abilities.
@@
Most existing open-source MLLMs perform poorly in the Multimodal Knowledge-Seeking (MKS) scenario. For instance, the performance of mPLUG-Owl2 in both categories of Vision-linked Textual Knowledge (VTK) and Text-linked Visual Knowledge (TVK) is below 30%. In the Fine-grained Visual Recognition (FVR) task, which requires the combination of fine-grained perception and comparison abilities, mPLUG-Owl2’s performance is merely over 10%. Compared to single-image MLLMs, multi-image models such as IDEFICS2-I show better capabilities in utilizing multimodal external knowledge. However, there is still significant room for improvement, as the performance of IDEFICS2-I on both VTK and TVK tasks is under 40%.
4.2.2 Multimodal In-Context Learning
[…]
4.3 Analysis
@@ 5 Conclusion
While MLLMs have shown strong performance in various vision-language tasks, their abilities with multi-image inputs remain underexplored. To address this gap, we introduce MIBench in this paper, a benchmark that evaluates MLLMs in three multi-image scenarios: multi-image instruction, multimodal knowledge-seeking, and multimodal in-context learning, across 13 tasks with 13K annotated samples. The evaluation results reveal that existing models, despite excelling in single-image tasks, face significant challenges with multi-image inputs. We will release our benchmark’s annotated data to facilitate further research. We hope this work spurs progress in improving the multi-image abilities of MLLMs.
6 Limitations
Due to the input length limitations of current MLLMs, our constructed benchmark includes 2 to 8 input images in each sample. However, real-world scenarios may involve a large number of images. We’ll investigate the evaluation with more images in future work.

### v2

@@ 4.2.1 Multi-Image Instruction & Multimodal Knowledge-Seeking
Closed-source MLLMs exhibit superior performance than open-source models. In most MII and MKS tasks, the performance of open-source models lags significantly behind that of proprietary models. For instance, on the Temporal Reasoning (TR) task, the best-performing open-source model MMICL achieves an accuracy of 47.0%, falling behind GPT-4o by 21.0%.
Open-source MLLMs are inadequate in fine-grained perception tasks. Although many open-source MLLMs have decent performance on the General Comparison (GC) task, their performance is significantly worse on the Subtle Difference (SD) and Text-Rich Images (TRI) VQA tasks. For instance, Idefics2-I achieves 83.1%, 49.7% and 43.9% on the three tasks respectively. In contrast, GPT-4V and GPT-4o largely outperform open-source models, due to their high-resolution input strategy. Figure 3 provides a qualitative case supporting the above point.
Multi-image MLLMs perform better than single-image models on most tasks. This verifies that pre-training on interleaved image-text data (e.g. Idefics2-I) and instruction tuning on multi-image data (e.g. Mantis) are both beneficial for improving multi-image abilities. Combining multi-image pre-training and instruction tuning, mPLUG-Owl3 achieves better performance than other open-source MLLMs on most tasks.
@@
Most existing open-source MLLMs perform poorly in the Multimodal Knowledge-Seeking (MKS) scenario. For instance, the accuracy rates of mPLUG-Owl2 on both the Vision-linked Textual Knowledge (VTK) and Text-linked Visual Knowledge (TVK) tasks are below 30%. In the Fine-grained Visual Recognition (FVR) task, which requires the combination of fine-grained perception and comparison abilities, mPLUG-Owl2’s performance is merely over 10%. Compared to single-image MLLMs, multi-image models such as Idefics2-I exhibit better capabilities in utilizing multimodal external knowledge. However, there is still significant room for improvement, as the performance of Idefics2-I on both the VTK and TVK tasks is under 40%.
@@ 5 Conclusion
While MLLMs have shown strong performance in various vision-language tasks, their abilities with multi-image inputs remain underexplored. To address this, we introduce MIBench in this paper, a benchmark that evaluates MLLMs across three multi-image scenarios: multi-image instruction, multimodal knowledge-seeking and multimodal in-context learning, covering 13 tasks with 13K annotated samples. The evaluation results reveal that existing models, despite excelling in single-image tasks, face significant challenges with multi-image inputs. The annotated data is publicly available to facilitate further research. We hope this work can spur progress in improving the multi-image abilities of MLLMs.
Limitations
Due to the input length limitation of current MLLMs, the Multi-Image Instruction and Multimodal Knowledge-Seeking scenarios of our benchmark include 2 to 8 input images in each sample. However, real-world scenarios may involve a large number of images. We’ll investigate the evaluation of MLLMs over more images in future work.
Acknowledgement
This work is supported by Beijing Natural Science Foundation (JQ21017, L243015, L223003), the National Key Research and Development Program of China (No. 2020AAA0105802), the Natural Science Foundation of China (No. 62036011, 62192782), and the Project of Beijing Science and Technology Committee (No. Z231100005923046).
Appendix A More Details of MIBench
Table 5 presents the detailed data generation information of each task. Note that “sampled from annotations” isn’t simple random sampling from the original annotations. Instead, as stated in Section 3.2, it involves designing specific sampling strategies tailored to the task.
Table 6 shows the detailed statistics of each task, including image number per sample, average question length and average answer length. Note that “Image Number Per Sample” for the Multimodal In-Context (MIC) learning scenario is a range determined by the number of demos. And the whole benchmark has 125K images in total. “Average Answer Length” refers to the average length of options for multiple-choice questions and the average length of answers for short-answer questions.